Q= how many possible committees can be formed from a group of seven people if each committee can have any number of member from 1 to 7?
the answer given is 127, solution pls

Answer by stanbon(75887)

how many possible committees can be formed from a group of seven people if each committee can have any number of member from 1 to 7?
the answer given is 127, solution pls
------------------
Ans:: # of subsets = 2^7 = 128
One of those sets is the empty set.
---
Answer to your question = 128-1 = 127
----------------------
Cheers,
Stan H.
----------------